OpenStudy (anonymous):

Rewrite y = x^2 + 14x + 29 in the form y = a(x - h)^2 + k.

OpenStudy (anonymous):

(x + ?)^2 = (x^2 + 2*x*? + ?^2)...........

OpenStudy (anonymous):

(u+v)^2 = u^2+2uv+v^2 in your case u=x but what is b? i give you a hint, solve: 2ab=14x when you have b you can find the missing k in your equation by calculating b^2-k then write down your new equation and look if you even need an a